OpenStudy (anonymous):

Natural Log (ln(x)) - When integrating a function, and the answer is ln(x), do we take the absolute value of x? example I have a Definite* Integral (3, 4) of [(1/(x-8))dx] - Definite Integral (3,4) of [(1/(x+9)]dx The answer, I know, comes out to be ln(x-8) - ln(x+9). so using the law of log, I can make that ln((x-8)/(x+9) What do I do with the negative values? is the antiderivative supposed to be the absolute value of ln(x) *i.e. ln|x|*?

OpenStudy (anonymous):

so i have ln(-48/-65)?

OpenStudy (dumbcow):

yes put the abs value inside log to take care of neg values

OpenStudy (dumbcow):

so yes answer is ln(48/65)

OpenStudy (anonymous):

does ln(48/65) = -ln(65/46)? @dumbcow

OpenStudy (dumbcow):

yes because of log property \[\ln(x^{-1}) = -\ln(x)\]

OpenStudy (anonymous):

okay, great.. i fully understand it now. thanks!